Status Report

ISS On-Orbit Status 7 Dec 2002

By SpaceRef Editor
December 7, 2002
Filed under , ,
ISS On-Orbit Status 7 Dec 2002

All ISS systems continue to function nominally, except as noted previously or below.  Today 30 years ago Apollo 17 was launched to the Moon, carrying Gene Cernan, Ron Evans and Harrison Schmitt on the last lunar landing mission of the Apollo Program.

The ISS crew enjoys a typical Saturday, off duty with only a few activities scheduled.

After wake-up at 4:00am EST, CDR Ken Bowersox and FE-1 Nikolai Budarin continued the renal (kidney) stone prevention experiment.  Bowersox had his second day of the regimen, which required him to collect urine samples in the course of the day besides diet logging.  For Budarin, sample collection ended this morning.

Afternoon activities consisted mostly of the weekly housecleaning by all crewmembers, today scheduled for 1.5 hours.  The periodic cleanup focuses on removal of food waste products, cleaning of compartments with vacuum cleaner, and wet cleaning of surfaces.

Time was again scheduled for remaining post-11A stowage operations by Bowersox.

The crew performed their regular physical exercise on TVIS (treadmill), CEVIS (cycle), RED (resistive) and VELO (cycle) with load trainer.

For FE-2/SO Don Pettit, a private family conference (PFC) was arranged for later today (3:45 pm) via S-band.

Pettit’s successful work yesterday with the HRF RIC (human research facility/rack interface controller) software load was gratefully acknowledged by POC (Payload Operations Center).  The rack will be activated next week for PuFF (pulmonary function in flight) activities.

On its fourth day of waved-off landing attempts (EOM+3), STS-113/Endeavour touched down at KSC on Orbit 215 at 2:37pm EST after a 5.7 million miles trip, concluding a highly successful ISS-11A mission.  Total duration: 13d 18h 47m 25s.  For the Expedition 5 crew of CDR Valery Korzun, FE-1/SO Peggy Whitson and FE-2 Sergei Treschev, this marks the 185th day of their stay in space (178 days aboard ISS).

Science Update (Expedition Six — 1st):

Extra-Vehicular Activity Radiation Monitors (EVARM):  In progress.

GASMAP/Pulmonary Function in Flight (PuFF):  PuFF is awaiting next week’s pre- and post-EVA sessions. Staff will be on console both days if any questions arise.  

Renal (Kidney) Stone Experiment:  Investigators hope that the first renal stone sample collections went well and are looking forward to getting the BCR (barcode reader) data from the crew’s diet and sample logging

Human Research Facility/Workstation (HRF WS):   The crew was thanked for successfully updating the RIC (rack interface controller) software, which clears the path for next week’s PuFF and FOOT experiments.  Attempts are underway to replan the Workstation checkout for the week of 12/9 in preparation for FOOT.

Solidification Using a Baffle in Sealed Ampoules (SUBSA):  Complete.  To be returned at next opportunity.

Pore Formation and Mobility Investigation (PFMI):  Complete for Inc. 5.  Remaining samples (7-8) to be processed when MSG (microgravity science glovebox) is available again.

Space Acceleration Measurement System (SAMS):  SAMS continues to collect acceleration data for vibratory characterization of the microgravity environment.

Microgravity Acceleration Measurement System (MAMS):  MAMS continues measurement of microgravity environment in the quasi -steady regime for general characterization.  

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):   PCG STES temperatures are nominal.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Zeolite Crystal Growth (ZCG):  Samples returned on 11A. New experiments planned.

Crew Earth Observation (CEO):  The ground received and reviewed downlinked DCS (digital camera system) imagery of the eclipse earlier this week over the southern hemisphere.  It was identified it in 23 frames and released to the public.  Some of them will be posted on the website.  Appreciation was expressed for the crew’s effort and dedication to acquire useful photos of this phenomenon under less than ideal conditions.

Today’s CEO targets were Angolan Biomass Burning (the fire season is coming to an end in southern Africa now.  However this region is a well illuminated on this pass and both fires and fire scars should have been easy to spot.  As ISS crossed northern Namibia from the SW, the crew was to try for oblique views to the left of track), Patagonian Glaciers (a brief break in the weather is expected for southern Patagonia for the next day or so. On this southernmost pass in low light, the crew was to look left of track in low-oblique or oblique views for context views of the Southern Patagonian Ice Field with its terrain enhanced by shadows. A second pass had better lighting to document details of the ice fields and glaciers, especially on the Pacific side if clouds permit. Crew was to use the long lenses of the DCS for nadir or near-nadir views), Lake Nasser, Toshka Lakes; Egypt (this expanding system of lakes to the NW of Lake Nasser is of keen interest for economic and environmental reasons. Egypt hopes to relocate its exploding population to the new arable lands around these lakes, but can this continued diversion of Nile River waters be sustained?  The ISS pass was particularly good for documenting the location and extent of these new water bodies with the possibility of sun glint enhancement.  Crew was to try for nadir or near-nadir views), Tuamotu Archipelago (although there may more clouds than usual for this target area, lighting is ideal to map the details of the coral reef structures of this large archipelago. ISS tracked over the widest and most dense portion, so the crew was asked to use the long lens of the DCS  for nadir views where clouds permit).

CEO images can be viewed at the website

SpaceRef staff editor.